What would occur to the total revenue coming from this product is you increase price to $11 per unit?
1
Expert's answer
2018-09-14T09:57:09-0400
The Revenue (R) is equal to the price (P) multiplied by the quantity of goods or services (Q).
R = P*Q.
The previous Revenue:
R1 = P1*Q1.
The Revenue after increasing of price:
R2 = P2*Q2.
We know: P2 = P1 + 11, Q2= *Q1
So, R2 = (P1 + 11)*Q1
As we see, the revenue would increase by R2 - R1 = (P1 + 11)*Q1 - P1*Q1 = 11Q1
Answer
The total revenue would increase by 11Q1, where Q1 is the number of units.